[figure: 9.7 cm]

SU(2) COHERENT STATE PATH INTEGRAL INVESTIGATION

OF THE HOLSTEIN DIMER

T. Hakioǧlu^{a}, V.A. Ivanov^{b}, and M.Ye. Zhuravlev^{b} aPhysics Department, Bilkent University, 06533 Ankara, Turkey bInstitute of General & Inorganic Chemistry of RAS, 31 Leninskii prosp., Moscow, 117907, Russia

The SU(2) coherent state path integral is used to investigate the partition function of the Holstein dimer. The SU(2) coherent state path integral takes naturally into account the dynamical symmetry of the model. Ground-state energy, the number of phonons and the average value of phonon coordinate are calculated as functions of parameters of the Hamiltonian. Renormalizations of the phonon frequency and electron hopping for bounding and antibounding states are considered. Comparison of the results obtained with the adiabatic and nonadiabatic limits is performed.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M.Ye. Zhuravlev

Address for correspondence : Institute of General & Inorganic Chemistry of RAS, 31 Leninskii prosp., Moscow, 117907, Russia

Email address : zhur@ionchran.rinet.ru